2) Jack owns a pizza place, where he works as the pizza chef. He pays himself a salary of $1,500. Jack has calculated that one pizza requires $0.50 worth of flour, $0.05 worth of yeast, $0.01 worth of water, $3.00 worth of cheese, and $2.00 worth of pepperoni to make. Jack also pays $3,000 for rent, $200 for insurance, $450 for utilities, and $500 for advertising. If Jack sells one pizza for $10, how many pizzas does he need to sell in a month to break even?
Answer #1

Selling price per unit = $10

variable cost per unit =  flour + yeast + water + cheese + pepperoni

= 0.50 + 0.05 + 0.01 + 3 + 2

= $5.56

Fixed cost = Salary + Rent + Insurance + Utilities + Advertising

= 1,500 + 3,000 + 200 + 450 + 500

= $5,650

Contribution margin per unit = Selling price per unit - variable cost per unit

= 10 - 5.56

= $4.44

Break even point in units = Fixed cost/Contribution margin per unit

= 5,650/4.44

= 1,272.52

= 1,273 (Rounded to next whole number)

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.